OpenStudy (anonymous):

Auto GOLD MEDAL for you if you answer Pro maths people come. binomial theorem: (3x-5)^4n-3 has 26 terms find the value of n

OpenStudy (anonymous):

i believe it should be 4n-3=26-1 n=7 because if you notice, a binomial expands out to one more term that its power

OpenStudy (anonymous):

@Auxuris hmm so why would you do 26-1? Is it because of the fact that \[T _{r+1}\] so it would be 4n-3=25? 25-3 = 22 22/4 = 5.5?... You plug in or no? Could you show a step by step sequence? :0 With the formulas and all?

OpenStudy (anonymous):

you mean 25+3...

OpenStudy (anonymous):

oh sure okay yes i used the general term method and got this 4n-3=26-1 4n=26-1+3 4n=28 n=7 when you plug it back in, (3x-5)^4n-3 = (3x-5)^4(7)-3 = (3x-5)^25 and if you expand it, you'll get 26 terms

OpenStudy (anonymous):

...oh.. sooo it would be 28/4 = 7 ahh gotcha on that part! Ahh gosh thanks so much!

OpenStudy (anonymous):

haha no prob ^^

OpenStudy (anonymous):

@Auxuris however, can I ask you another question, i'll give you another medal :D I'm struggling on how to answer this: Find c given that the expansion (1+cx)(1+x)^4 includes the term 22x^3 I know you need 4 choosing 3 something, with nCr on the calculator, how would I actually do this?

OpenStudy (anonymous):

you can't give two medals xD but it's okay, i don't come here for medals excellent, you're giving me practice on something i need so you need to compare coefficients say look at the right side, (1+x)^4 you're right, its a 4C3, which when expanded looks like  4x^3 will then be paired with the 1 in (1+x) in order to get something with x^3 follow? (: it would be, (4x^3)(1)=4x^3